sqrt(t+12)-t=0 is the problem. solve algebraically and check potential solutions
I'm not sure whether to add t+12 then square or distribute after t^2+12^2-t=0
: sqrt(t+12)-t=0 is the problem. solve algebraically and check potential solutions
I'm not sure whether to add t+12 then square or distribute after t^2+12^2-t=0

Answer by checkley77(3412)

sqrt(t+12)-t=0
SQRT(T+12)=T
NOW SQUARE BOTH SIDES.
(T+12)=T^2
T^2-T-12=0
(T-4)(T+3)=0
T-4=0 T=4 ANSWER.
T=3=0
T=-3 ANSWER.
Question 151453: sqrt(t+12)-t=0 is the problem. solve algebraically and check potential solutions
I'm not sure whether to add t+12 then square or distribute after t^2+12^2-t=0
: sqrt(t+12)-t=0 is the problem. solve algebraically and check potential solutions
I'm not sure whether to add t+12 then square or distribute after t^2+12^2-t=0

Answer by oscargut(667)

sqrt(t+12)-t=0
then
sqrt(t+12)=t
then
t+12=t^2
t^2-t-12=0
then
t= (1+7)/2 or t= (1-7)/2
then t=4 or t=-3 (potential solutions)
if t=4
sqrt(4+12)-4=sqrt(16)-4=0 (4 is a solution)
if t=-3
sqrt(-3+12)-(-3)=sqrt(9)+3=6 (-3 is not a solution)
Solution : t=4